# Hiba Arnaout1, Noy Sternlicht2, Tom Hope2,3, Iryna Gurevych1

### v1

@@ 3.3 Ablation study for impact summary generation
Prompt variants. We experiment with different input configurations to assess their effect on the summary’s quality, focusing on: (1) Citation contexts, including all, only impact-revealing, or none, to test what the LLM infers without citations and how varying context types affect the summary. Although full citation contexts can enrich summaries, they can also introduce noise from incidental references Liu et al. (2023a). (2) Fine-grained intents, whether to include the intents alongside citation contexts.
Table 3 lists all variants. In total, we generate 945 impact summaries6661 variant without citations + (4 citation-based variants * 2 orderings) * 105 papers. See Appendix C for the prompt and schema. For the intent generation task, we choose GPT-4o-mini due to its lower cost. For both impact generation and evaluation, the more advanced tasks, we use GPT-4o.
@@
Results. Table 3 shows the results for all metrics. Even though it shows a good understanding of how the paper was cited, likely due to the LLM’s training data having access to a large corpus of scholarly articles, the zero-shot variant is the least faithful, and we observe that prompts with only impact-revealing citations generate more faithful impact summaries. This suggests that longer contexts might induce hallucinations in the generation process. Interestingly, providing the citation intents also gives a slight boost to the faithfulness of the impact summaries. A possible reason is that the intents encourage wording that aligns with the input paper’s citations.
@@
Coverage is especially challenging, requiring a balance of completeness, relevance, and conciseness. Still, our variant outperforms the baseline by 9%, showing the LLM can capture key themes. Some inputs contain over 100 distinct themes, explaining modest overall scores. When evaluating only the top-3 most frequent themes (the largest 3 clusters), coverage rises significantly, as our best variant reaches 0.75 (+17%), highlighting its ability to capture the most important aspects of a paper’s impact. Variants exposed only to impact-revealing citations show higher citation year compliance, while adding citation intents lowers it, likely because models focus more on intent than timing. Examples are in Appendix G.1, with further discussion in Section 7.
All variants score high on trend awareness, with a slight edge for the impact-revealing + intents variant, showing LLMs can recognize distinct impact periods. This variant also improves insightfulness and specificity by 13% over the zero-shot baseline. Qualitative examples are in Appendix G.1. G-Eval, via DeepEval [^7] https://github.com/confident-ai/deepeval, also offers reasons for each score (examples in Appendix G.4). We visualize sample impact summaries in Appendix G.5. To examine whether results change depending on the field of the paper, we report results per field in Appendix G.2.
LLM-Human correlation We compute F1 agreement between human and LLM evaluators (50 instances per metric), yielding scores of 0.83 for faithfulness, 0.71 for coverage, 0.66 for insightfulness, 0.86 for trend awareness, and 0.74 for specificity. Correlation metrics (Spearman and Kendall-Tau) in Appendix G.3 show moderate to strong statistical significance.
@@ 3.4 Practical applications
We showcase the usability of our work in two practical applications. First, by analyzing 10 diverse research topics, such as LLMs for code generation, open information extraction, commonsense knowledge mining, and comparing papers with similar citation counts, it shows that papers can have very different types of impact. For example, in open information extraction, some papers are cited to highlight limitations or motivate new work Cui et al. (2018), while others are cited mainly for their methods Gashteovski et al. (2017); Han et al. (2019). More details and examples are in Appendix I.1.
Second, we explore generating author-level impact summaries by aggregating LLM-generated paper-level summaries for an author’s top-cited papers. Examples for two senior NLP researchers are shown in Appendix I.2. This approach has strong potential for broader applications at the level of labs, institutions, or venues, which we leave for future work.
@@ 4 Human Evaluation
Setup. To gain deeper insights into the quality of our impact summaries and collect expert feedback for potential future improvements, we conduct a user study with 9 university professors from diverse backgrounds (details in Appendix H.1). Each expert reviews impact summaries of their own papers through a two-part evaluation: (1) a pairwise comparison, where they choose the better summary based on relevance and insightfulness, comparing the zero-shot variant against our best variant. To prevent bias, we alternate the positions of the impact summaries. (2) a perceived usefulness assessment, where experts rate their agreement with given statements about an impact summary (generated by our best variant), using a 1-5 Likert scale. The professors evaluated the impact summaries for 82 papers and offered open-ended feedback.
Results. Figure 3 shows results of the pairwise comparison. Impact summaries generated by our variant [^8] Impact-revealing citations with intents. demonstrate greater relevance (63%) and insightfulness (75%) than the baseline [^9] Zero-shot variant.
Results of the perceived usefulness (visualized in Appendix H.2) show that approximately 60% of professors agreed that the summaries provided an appropriate level of detail (clarity) and offered novel insights into how their papers were used, i.e., information not readily available elsewhere. Notably, for papers in the top 10% based on impact-revealing citations, agreement on clarity and informativeness rose to 75%. Results and a discussion on open feedback are in Appendix H.3.
@@ 5 Related Work
Scientific impact analysis. Studying scientific impact is vital for guiding research and recognizing contributions. While prior work relies on citation counts and related metrics Hutchins et al. (2016); Bos and Nitza (2019); Siudem et al. (2020); Min et al. (2021); Wahle et al. (2023); Gu and Krenn (2024), impact cannot be reduced to a single number Zhu et al. (2015). Few studies use citation context to assess scientific impact, and those that do focus on classifying citation intent individually rather than aggregating them in summaries Jurgens et al. (2018); Valenzuela et al. (2015). Our approach uses fine-grained citation context analysis to capture and aggregate specific citation reasons across impact-revealing citations.
Other work links impact to novelty of the paper’s content Rüdiger et al. (2021); Shi and Evans (2023); Arts et al. (2021), but overlooks how papers are used. Most also equate impact with praise Zhang et al. (2024); Valenzuela et al. (2015); Zhu et al. (2015), whereas we distinguish confirmatory from correction citations, offering a fuller view. As detailed in Appendix A, we are the first to generate time-aware, multifaceted impact summaries from fine-grained citation analysis.
@@
Citation intent prediction. Prior work on citation intent classification uses pre-defined categories (e.g., method use, background) with early methods relying on supervised learning and linguistic features Teufel et al. (2006); Jurgens et al. (2018); Tuarob et al. (2019). Later approaches applied multi-task learning Cohan et al. (2019), while Lauscher et al. (2022) introduced a multi-label method using fine-tuned SciBERT Beltagy et al. (2019). These rely on coarse categories; in contrast, we use in-context learning to generate fine-grained, free-form intent descriptions (Table F.2), offering more precise insights, an approach recently shown to aid related work generation Şahinuç et al. (2024).
@@
Multi-document summarization in the era of LLMs. Recent work in query-focused multi-document summarization uses LLMs to generate context-aware summaries Roy and Kundu (2023), with approaches like graph-based QA over private corpora Edge et al. (2024) and retrieval-augmented summarization Zhang et al. (2025). In the scientific domain, prior efforts focus on summarizing papers for related work Şahinuç et al. (2024); Lu et al. (2020); Chen et al. (2021); Wu et al. (2021) or reviews Kasanishi et al. (2023). In contrast, we generate free-text summaries of scientific impact using a novel time-aware, multi-document approach.
@@
LLM Evaluators. Retrieval-Augmented Generation (RAG) systems condition generation on external information retrieved from user queries Yu et al. (2024); Gao et al. (2023), but evaluating them is challenging due to multiple stages like chunking and retrieval. Tools like RagChecker Ru et al. (2024) use entailment-based metrics to assess faithfulness, while RAGAs Shahul et al. (2023) offers reference-free evaluation of relevance and faithfulness. Inspired by these, we assess our impact summaries’ faithfulness and coverage by validating them against the citation context corpus.
The LLM-as-a-judge paradigm, used for tasks like QA and writing evaluation Zheng et al. (2023); Shao et al. (2024); Asai et al. (2024), enables scalable assessment of complex qualities. Given the novelty of scientific impact summarization, we adopt G-Eval Liu et al. (2023b), a CoT-based framework, to assess informativeness.
@@ 6 Conclusion
This work introduces a novel approach for generating time-aware impact summaries of research papers by analyzing evolving fine-grained confirmatory and correction citation intents. Our approach goes beyond citation counts, offering a more nuanced understanding of a paper’s impact over time. Automated evaluations align moderately to strongly with human judgments, and expert feedback highlights the value and potential for refinement of our summaries. We release our data and code to support research in this new task.
@@ 7 Limitations
Trustworthiness evaluation. Without restrictions, our approach shows unsatisfying coverage (0.34) and citation year compliance (0.59), compared to other metrics such as faithfulness. Regarding the year compliance, citation phrases often include multiple references accompanied by publication years, which can confuse the model and lead to incorrect citation year associations. In the future, we plan to remove or distinguish unrelated, potentially distracting numbers from the citation context (e.g., using well-crafted heuristics or prompt-based methods). For coverage, the LLM may focus on certain prominent themes while underrepresenting others, especially with complex topics. We briefly reported the coverage numbers for the most frequent themes, which showed a great improvement. In the future, we plan to explore this further, for instance, by ranking the importance of the themes and factor this into the coverage computation at different ranks. However, for this study, we decided to test full coverage to make sure the model captures a wide range of themes, without focusing on any specific ones, to check its overall understanding of the paper’s impact. Finally, although checking potentially hundreds of citation contexts/intents at once raises concerns, the strong LLM-human agreement reported in Section 3.3, 0.83 F1 agreement for faithfulness and 0.71 for coverage, suggests such issues are unlikely.
@@ Appendix A Existing Work on Scientific Impact Analysis
Zhang et al. (2024) Influential-citation counts P. LLM Computer science
@@ Appendix C Prompt for Generating Impact Summaries and Output Schema
Generating an impact summary about a research paper The scientific impact summary of a research paper describes the impact a given paper had on other papers, including both praise and critique. To understand the impact of a paper, one needs to understand how exactly it has been utilized and discussed by other papers. This is normally referred to as citation intents. One also needs to understand the evolution of the impact and citation intents over time. Given an input paper’s title, its publication year, and its citation context, describe the impact of that paper. The citation context includes five components: $<$citation ID, citation title, citation year, citation context, citation intent$>$. Given the input paper with id \$paperId\$ titled \$title\$ published in \$year\$, and the following list of papers citing it: \$citation_context\$ Generate an impact summary about the input paper.
@@ E.2 More on setup and results
Figure 10 shows the quantitative results per paper field, and Table 7 shows qualitative examples. We do not observe any large variations indicating that our method is not limited to a specific field. However, the highest recall (a perfect score of 1.0) was achieved in the psychology field. This may suggest that citation contexts in psychology papers tend to express intentions more explicitly, particularly those that reveal impact, e.g. “Some of those assumptions have been controversial, as researchers disagree about whether the kinds of behaviors measured by particular implicit tests should be considered indicators of attitudes or something else”. This seems to be in line with our observations made during error analysis. To understand where our ICL prompt failed, we plot the confusion matrices in Figure 11 and inspect the 10% impact-revealing citation phrases that were misclassified as “other”. We notice that the classifier considers minor “resource use” intent as non-impact revealing, e.g., “We retrieved 400 results for each keyword, resulting in a total of 145,682 stories downloaded in the JavaScript Object Notation (JSON) format [48].” On the same note, it appears that when “inspiration” is only hinted at in the citation context, the classifier is not able to pick up the signal, e.g., “The search problem here corresponds closely to a binary dynamic constraint satisfaction problem [Mittal & Falkenhainer , 1990]”. Inter-annotator agreement on the annotated results is in Table 6. They range from good to strong.
@@
Task F1 agreement
Impact-revealing classifier 0.86
Fine-grained intent: correctness 0.77
Fine-grained intent: conciseness 0.85
@@
Appendix F Intents across Fields
In Table 2, we show an analysis of citation intents in the psychology, medicine, and computer science fields, each also divided by recency and citation counts. It shows that, generally, psychology citations tend to lean towards impact-revealing, computer science towards the non-impact-revealing (“other”) side, with the exception of more recent papers, and medicine has a more balanced split. Looking at specific examples, we notice that citations in psychology papers lack an objective tone, especially when it comes to critique, e.g., “some of those assumptions have been controversial”, and “researchers disagree about whether the kinds of behaviors measured by particular implicit tests should be considered indicators of attitudes or something else”. However, it does seem that in computer science there is a noticeable shift toward more impact-revealing citations in recent years. This shift likely reflects the novelty of recent research and the immediate impact of AI papers, as these papers have not yet entered the “legacy” phase, i.e., when papers start to receive more citations with “background work”-like intents. It is worth mentioning that these “legacy” papers are naturally strongly present in the highly cited subsets (by citation counts).

### v2

@@ 3.3 Ablation study for impact summary generation
Prompt variants. We experiment with different inputs to assess their effect on the summary’s quality, focusing on: (1) Citation contexts: including all, only impact-revealing, or none, to test what the LLM infers without citations and how varying context types affect the summary. Although full citation contexts can enrich summaries, they can also introduce noise from incidental references Liu et al. (2024). (2) Intents: whether to include the intents alongside citation contexts. Table 5 lists all variants. In total, we generate 945 impact summaries6661 variant without citations + (4 citation-based variants * 2 orderings) * 105 papers. Appendix C shows the prompt. For intent generation, we choose GPT-4o-mini due to its lower cost. For both impact generation and evaluation, the more advanced tasks, we use GPT-4o.
@@
Results. Table 5 shows the results for all metrics. Even though it shows a good understanding of how the paper was cited, likely due to the LLM’s training data having access to a large corpus of scholarly articles, the baseline (no-knowledge) variant is the least faithful, and we observe that prompts with only impact-revealing citations generate more faithful summaries. This suggests that longer contexts might induce hallucinations in the generation process. Interestingly, providing the citation intents also gives a slight boost to the faithfulness of the summaries, possibly because the intents encourage wording that aligns with the input paper’s citations.
@@
Coverage is especially challenging, requiring a balance of completeness, relevance, and conciseness. Our best variant still outperforms the baseline by 9%, showing that the LLM can capture key themes. Some inputs include over 100 distinct themes, explaining modest overall scores. Focusing only on the 3 most frequent themes (largest clusters), coverage rises significantly, as our best variant reaches 0.75 (+17%), highlighting its strength in capturing core impact. To assess the coverage of meaningful low-frequency themes, we analyze long-tail impact across 10 papers; Appendix G.2 shows up to 50% coverage when intents are included. Variants using only impact-revealing citations show higher citation year compliance; adding citation intents lowers it, likely due to greater focus on intent over timing. See Appendix G.1 for examples and Section 7 for discussion. All variants show high trend awareness, with a slight edge for the impact-revealing + intents variant, showing LLMs can recognize distinct impact periods. This variant also improves insightfulness and specificity by 13% over the no-knowledge baseline. Qualitative examples are in Appendix G.1. G-Eval, via DeepEval [^7] https://github.com/confident-ai/deepeval, also offers reasons for each score (Appendix G.5). We visualize sample summaries in Appendix G.6. To examine whether results change depending on the paper’s field, we report results per field in Appendix G.3. Finally, Appendix G.7 show insights into how intent misclassification can impact the quality of the generated summaries.
Beyond GPTs. We also generate impact summaries using Qwen and Gemini, to ensure model-agnosticism. Quantitative and qualitative results are provided in Appendix G.8. Although GPT-4o served as our primary model, both Qwen and Gemini demonstrated strong performances, with Gemini in particular standing out in citation year compliance (0.93) and faithfulness (0.96).
LLM-Human correlation We compute the human-LLM correlation, which shows moderate to strong relationships for the Spearman and Kendall-Tau metrics. Details are provided in Appendix G.4.
@@ 3.4 Practical applications
We showcase the usability of our work in two practical applications. First, by analyzing 10 diverse research topics, such as LLMs for code generation, open information extraction, commonsense knowledge mining, and comparing papers with similar citation counts, it shows that papers can have very different types of impact. For example, in open information extraction, some papers are cited to highlight limitations or motivate new work Cui et al. (2018), while others are cited mainly for their methods Gashteovski et al. (2017); Han et al. (2019). More details and examples are in Appendix I.1. Second, we explore generating author-level impact summaries by aggregating LLM-generated paper-level summaries for an author’s top-cited papers. Examples for two senior NLP researchers are shown in Appendix I.2. This method could be extended to institutions and venues, which we leave for future work.
@@
Comprehensive details of the models and their configurations for each task/component are provided in Appendix H to support reproducibility and attribution.
@@ 4 Human Evaluation
Setup. To gain deeper insights into the quality of our impact summaries and collect expert feedback for potential future improvements, we conduct a user study with 9 university professors from diverse backgrounds  [^8] Nationalities: 2 German, 2 British, 2 Chinese, 1 American, 1 Czech, 1 Albanian; genders: 5 male, 4 female; research areas: AI, NLP, knowledge graphs, databases and information systems, computational social sciences, psychology and brain sciences.. Each expert reviews impact summaries of their own papers through a two-part evaluation: (1) a pairwise comparison, where they choose the better summary based on relevance and insightfulness, comparing the no-knowledge variant against our best variant. To prevent bias, we alternate the positions of the impact summaries. (2) a perceived usefulness assessment, where experts rate their agreement with given statements about an impact summary (generated by our best variant), using a 1-5 Likert scale. The professors evaluated the impact summaries for 82 papers and offered open-ended feedback.
Results. Figure 2 shows results of the pairwise comparison. Impact summaries generated by our variant [^9] Impact-revealing citations with intents. demonstrate greater relevance (63%) and insightfulness (75%) than the baseline101010no-knowledge variant. Results of the perceived usefulness, in Figure 3, show that approximately 60% of professors agreed that the summaries provided an appropriate level of detail (clarity) and offered novel insights into how their papers were used, i.e., information not readily available elsewhere. Notably, for papers in the top 10% based on impact-revealing citations, agreement on clarity and informativeness rose to 75%. Finally, we ask the evaluators for open feedback. A couple professors found certain summaries too generic, failing to highlight specific strengths and limitations. They also pointed out that summaries based on citations from lesser-known conferences may not fully reflect a paper’s true influence, and the quality of summaries often correlates with the size of impact-revealing citation context (a point further validated in Figure 3 (b)). Additionally, one professor noted that for some papers, it was difficult to assess impact because the work was collaborative, and they were only familiar with certain aspects of the paper’s contributions, while their co-authors might be more familiar with other aspects. Other professors suggested improvements in both content and structure: they recommended incorporating citation counts within the text, which we plan to implement in the future, and enhancing the structure of our semi-structured summaries by adding elements like bullet points to improve readability.
Figure 3: We show the overall results in (a), and results for two subsets of papers, namely papers with the most impact-revealing citations in (b), and papers with the highest number of citations in (c). Both subsets show impact summaries that are rated higher, likely because higher counts of impact-revealing citations and citations in general allow for richer contexts, which provide the model with clearer signals about the paper’s impact, leading to more accurate and insightful summaries.
@@ 5 Related Work
Scientific impact analysis. Studying scientific impact is vital for guiding research and recognizing contributions. While prior work relies on citation counts and related metrics Hutchins et al. (2016); Bos and Nitza (2019); Siudem et al. (2020); Min et al. (2021); Wahle et al. (2023); Gu and Krenn (2024), impact cannot be reduced to a single number Zhu et al. (2015). Few studies use citation context to assess scientific impact, and those that do focus on classifying citation intent individually rather than aggregating them in summaries Jurgens et al. (2018); Valenzuela et al. (2015). Our approach uses fine-grained citation context analysis to capture and aggregate specific citation reasons across impact-revealing citations. Other work links impact to novelty of the paper’s content Rüdiger et al. (2021); Shi and Evans (2023); Arts et al. (2021), but overlooks how papers are used. Most also equate impact with praise Zhang et al. (2024a); Valenzuela et al. (2015); Zhu et al. (2015), whereas we distinguish confirmatory from correction citations, offering a fuller view. As detailed in Appendix A, we are the first to generate time-aware, multifaceted impact summaries from fine-grained citation analysis.
@@
Citation intent prediction. Prior work on citation intent classification uses pre-defined categories (e.g., method use, background) with early methods relying on supervised learning and linguistic features Teufel et al. (2006); Jurgens et al. (2018); Tuarob et al. (2020). Later approaches applied multi-task learning Cohan et al. (2019), while Lauscher et al. (2022) introduced a multi-label method using fine-tuned SciBERT Beltagy et al. (2019). These rely on coarse categories; in contrast, we use in-context learning to generate fine-grained, free-form intent descriptions (Table 11), offering more precise insights, an approach recently shown to aid related work generation Sahinuç et al. (2024).
@@
Multi-document summarization in the era of LLMs. Recent work in query-focused multi-document summarization uses LLMs to generate context-aware summaries Roy and Kundu (2024), with approaches like graph-based QA over private corpora Edge et al. (2024) and retrieval-augmented summarization Zhang et al. (2024b). In the scientific domain, prior efforts focus on summarizing papers for related work Sahinuç et al. (2024); Lu et al. (2020); Chen et al. (2021); Wu et al. (2021) or reviews Kasanishi et al. (2023). In contrast, we generate free-text summaries of scientific impact using a novel time-aware, multi-document approach.
@@
LLM Evaluators. Retrieval-Augmented Generation (RAG) systems condition generation on external information retrieved from user queries Yu et al. (2024); Gao et al. (2023), but evaluating them is challenging due to multiple stages like chunking and retrieval. Tools like RagChecker Ru et al. (2024) use entailment-based metrics to assess faithfulness, while RAGAs ES et al. (2024) offers reference-free evaluation of relevance and faithfulness. Inspired by these, we assess our impact summaries’ faithfulness and coverage by validating them against the citation context corpus. While inspired by RAG methods, our approach introduces a distinct task focused on impact and temporal dynamics. By targeting impact-revealing citations and modeling shifts over time, we support more nuanced, citation-centric summarization. The LLM-as-a-judge paradigm, used for tasks like QA and writing evaluation Zheng et al. (2023); Shao et al. (2024); Asai et al. (2024a), enables scalable assessment of complex qualities. Given the novelty of scientific impact summarization, we adopt G-Eval Liu et al. (2023b), a CoT-based framework, to assess informativeness.
@@ 6 Conclusion
This work introduces a novel approach for generating time-aware impact summaries of research papers by analyzing evolving fine-grained confirmatory and correction citation intents. Our approach goes beyond citation counts, offering a more nuanced understanding of a paper’s impact over time. Expert evaluation highlights the value and potential for refinement of our summaries. We release our data and code to support research in this new task.
@@ 7 Limitations
Trustworthiness evaluation. Without restrictions, our approach shows unsatisfying coverage (0.34) and citation year compliance (0.59), compared to other metrics such as faithfulness. Regarding the year compliance, citation phrases often include multiple references accompanied by publication years, which can confuse the model and lead to incorrect citation year associations. In the future, we plan to remove or distinguish unrelated, potentially distracting numbers from the citation context (e.g., using well-crafted heuristics or prompt-based methods). For coverage, the LLM may focus on certain prominent themes while underrepresenting others, especially with complex topics. We briefly reported the coverage numbers for the most frequent themes, which showed a great improvement. In the future, we plan to explore this further, for instance, by ranking the importance of the themes and factor this into the coverage computation at different ranks. However, for this study, we decided to test full coverage to make sure the model captures a wide range of themes, without focusing on any specific ones, to check its overall understanding of the paper’s impact. Finally, although checking potentially hundreds of citation contexts/intents at once raises concerns, the strong LLM-human agreement reported in Appendix G.4 suggests such issues are unlikely.
@@
Broader spectrum of scientific impact. While our study operationalizes scientific impact primarily through confirmation and correction, we acknowledge that real-world impact spans a broader spectrum (e.g., parallel development). Our framework is designed to be extensible, and future work can enrich the training examples with additional categories to capture a more nuanced landscape of scholarly influence, though we note that such categories may also introduce additional challenges and interpretive ambiguity.
Data reliability and quality controls. We rely on the S2AG resource from Semantic Scholar, whose scale, coverage, active curation, and provision of citation contexts and structured metadata make it a uniquely reliable foundation for studying citations. While the data is generally high quality, our pipeline adds safeguards such as removing duplicate contexts, discarding citations lacking sufficient information, and enforcing temporal consistency to ensure robustness. We acknowledge that very large-scale datasets may have completeness limitations, but such issues are common across bibliographic resources, and further citation-quality filtering is a natural future enhancement rather than a prerequisite for the validity of our current findings.
@@
Acknowledgments
This work has also been co-funded by the LOEWE Distinguished Chair “Ubiquitous Knowledge Processing”, LOEWE initiative, Hesse, Germany (Grant Number: LOEWE/4a//519/05/00.002(0002)/81 and by the European Union (ERC, InterText, 101054961). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Existing Work on Scientific Impact Analysis
Zhang et al. (2024a) Influential-citation counts P. LLM Computer science
@@ Appendix C Prompt for Generating Impact Summaries and Output Schema
Generating an impact summary about a research paper The scientific impact summary of a research paper describes the impact a given paper had on other papers, including both praise and critique. To understand the impact of a paper, one needs to understand how exactly it has been utilized and discussed by other papers. This is normally referred to as citation intents. One also needs to understand the evolution of the impact and citation intents over time. Given an input paper’s title, its publication year, and its citation context, describe the impact of that paper. The citation context includes five components: <<citation ID, citation title, citation year, citation context, citation intent>>. Given the input paper with id \$paperId\$ titled \$title\$ published in \$year\$, and the following list of papers citing it: \$citation_context\$ Generate an impact summary about the input paper.
@@ E.2 More on setup and results
Figure 10 shows the quantitative results per paper field, and Table 9 shows qualitative examples. We do not observe any large variations indicating that our method is not limited to a specific field. However, the highest recall (a perfect score of 1.0) was achieved in the psychology field. This may suggest that citation contexts in psychology papers tend to express intentions more explicitly, particularly those that reveal impact, e.g. “Some of those assumptions have been controversial, as researchers disagree about whether the kinds of behaviors measured by particular implicit tests should be considered indicators of attitudes or something else”. This seems to be in line with our observations made during error analysis. To understand where our ICL prompt failed, we plot the confusion matrices in Figure 11 and inspect the 10% impact-revealing citation phrases that were misclassified as “other”. We notice that the classifier considers minor “resource use” intent as non-impact revealing, e.g., “We retrieved 400 results for each keyword, resulting in a total of 145,682 stories downloaded in the JavaScript Object Notation (JSON) format [48].” On the same note, it appears that when “inspiration” is only hinted at in the citation context, the classifier is not able to pick up the signal, e.g., “The search problem here corresponds closely to a binary dynamic constraint satisfaction problem [Mittal & Falkenhainer , 1990]”. Inter-annotator agreement on the annotated results is in Table 8. They range from fair to substantial. In this evaluation, annotators also indicate their confidence level in their decision on the impact-revealing classification as either “low”, “medium”, or “high”. If there is a disagreement, the annotation with the higher confidence level takes precedence. If the confidence levels are tied, the tie is resolved randomly.
@@
Task Cohen’s kappa
Impact-revealing classifier 0.68
Fine-grained intent: correctness 0.20
Fine-grained intent: conciseness 0.34
@@
Appendix F Effect of Different Number of Shots (K)
We conduct an experiment to test our ICL prompt with different Ks. We manually polished the 200 instances from the previous experiment and randomly split the set into train 40%, dev 30%, test 30%. For the number of shots, we test from K=1 to 80. For every dev and test instance, we run the prompt 5 times and report the average, using GPT4o-mini, shuffling the order of the shots. Results in Figure 12 show that increasing the number of shots has an effect on all metrics, outperforming both the zero-shot, and the always-impact-revealing baselines. Performance improves sharply up to K = 50, after which gains are minimal. At K = 50, metrics plateau with precision at 0.90, recall at 0.94, F1 at 0.92, and accuracy at 0.92.
Figure 12: Effect of different number of shots (K) on detecting impact-revealing citations. Zero-shot: horizontal dashed lines; always impact-revealing: dotted lines. This shows that adding a reasonable number of examples (around 50) is sufficient to yield a 60% improvement in recall.